OpenStudy (kayla2001):

y=4x and y + x = 5 a.)(-4,1) b.)(1,4) c.)(-3,2) d.)(2,3)

syed98 (syedmohammed98):

its B :)

OpenStudy (anonymous):

x=5-y, y=4(5-y), y=20-4y, 5y=20, y=4

OpenStudy (anonymous):

substitute y to solve for x

OpenStudy (kayla2001):

is b right ?

OpenStudy (anonymous):

yea

syed98 (syedmohammed98):

:)

OpenStudy (kayla2001):

I got more yall and I give medals for the best answers

OpenStudy (anonymous):

y=4, x=5-y, x=5-(4), x=1

OpenStudy (kayla2001):

y=2x-1 and y = x +3

syed98 (syedmohammed98):



OpenStudy (anonymous):

x+3=2x-1 x-2x=-3-1 -x=-4 x=4

syed98 (syedmohammed98):

i hope u get it :)

OpenStudy (anonymous):

y=(4)+3 y=7